OpenStudy (anonymous):

transform this function in vertex form f(x)= (3x+1)(2x-3) ---> PLLEEEAASSEE HELP ME :(

OpenStudy (anonymous):

6\[x^2\]-9x+2x-3 6\[x^2\]-7x-3

OpenStudy (anonymous):

how can i form it in vertex form?

OpenStudy (anonymous):

from up

OpenStudy (anonymous):

yes, but i need the vertex form, that is not the final answer :(

OpenStudy (anonymous):

\[6x^{2}\]-7x-3

OpenStudy (anonymous):

yes that is the first step, i need the 2nd and 3rd step :(

OpenStudy (anonymous):

no, thats last step

OpenStudy (anonymous):

no, but thanks for helping me..

OpenStudy (anonymous):

i've write first and second steps

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(x) = 2(x2 - 4x) + 7 (factor out 2 from the terms 2x2 - 8x) = 2(x2 - 4x + 4) + 7 - 8 (complete the square of x2 - 4x) = 2(x - 2)2 - 1 (factor the perfect square and simplify.)

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

but i dont get it

OpenStudy (anonymous):

first JHS?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

hmm thats 1grade JHS lesson

OpenStudy (anonymous):

nope, its SHS

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

i can finish it, but i don't know how to write it

OpenStudy (anonymous):

aw :(...huhu pls help me..:(, its a homework for tomorrow

OpenStudy (anonymous):

yeah, and i have homework for tomorrow too

OpenStudy (anonymous):

aw, sorry for disturbing you, hehe nice meeting you, thanks for effort :). THANK YOU! :))

OpenStudy (anonymous):

haha okay nope

OpenStudy (anonymous):

^________________________^

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."